OpenStudy (anonymous):

please help me 1. A red hot air balloon is 10 meters above the ground, rising 15 meters per minute, and a blue balloon is 150 meters above the ground, descending 20 meters per minute. How high in the air will they be when they pass each other? a. 60 meters b. 70 meters c. 4 minutes d. 90 meters 2. When solving the following system, which expression could be substituted for x? x + 4y = 1 2x - 3y = -9 a . 4y - 1 b. 1 - 4y c. 3y - 9 d. -9 - 3y 3. If x - 3y = -9 and 5x - 2y = 7, what is the value of x? a. 1 b. 2 c. 3 d. 4

OpenStudy (anonymous):

what do you think 1 is ?

OpenStudy (anonymous):

you have to help me as well xD

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

a

OpenStudy (anonymous):

b =)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

find the one they = to 1st xD

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

plz @jakashaka123 explain to crazy_girl how you got your answer

OpenStudy (welshfella):

don't just give the answer - Read Code of Conduct

OpenStudy (anonymous):

so you would add 15 to 10 and subtract 20 from 150

OpenStudy (anonymous):

ok she gots it

OpenStudy (anonymous):

eventually they will get to be the same number

OpenStudy (anonymous):

did i do good @welshfella

OpenStudy (anonymous):

hmmmm ????

OpenStudy (anonymous):

oh well my teacher has to take my phone thanks @welshfella

OpenStudy (anonymous):

can u help me with the rest

OpenStudy (welshfella):

yes that's better

OpenStudy (welshfella):

2. x + 4y = 1 make x the subject by subtracting 4y from both sides:- x = 1 - 4y

OpenStudy (welshfella):

so thats b for q 2

OpenStudy (anonymous):

thank you

OpenStudy (welshfella):

3. x - 3y = -9 5x - 2y = 7 we need to find the value of x best way is to multiply the first equation by 2 and second equation by -3 so we get 6y and -6y in the 2 equations 2x - 6y = -18 -15x + 6y = -21 Adding these 2 equations eliminate the term in y:- -13x = -39 divide both sides by -13:- x = 3

OpenStudy (anonymous):

thank u

OpenStudy (welshfella):

yw